OpenStudy (anonymous):

the expression 6 radical sign 20 divided by 3 radical sign 5 is equivalent to...

OpenStudy (saifoo.khan):

\[\frac{6 \sqrt{20}}{3 \sqrt5} \to 4\]

OpenStudy (anonymous):

its 4?

OpenStudy (anonymous):

how did u do it can u show me work pls

OpenStudy (anonymous):

2\sqrt{4}\] is what i get

OpenStudy (anonymous):

\[6 \sqrt20 / 3 \sqrt 5\]\[2 \sqrt20/\sqrt5\]\[2 \sqrt4\]\[2 \times 2 = 4\]

OpenStudy (saifoo.khan):

Yup!